# A domain decomposition-based autoregressive deep learning model for unsteady and nonlinear partial differential equations

### v1

@@ Abstract
In this paper, we propose a domain-decomposition-based deep learning (DL) framework, named transient-CoMLSim, for accurately modeling unsteady and nonlinear partial differential equations (PDEs). The framework consists of two key components: (a) a convolutional neural network (CNN)-based autoencoder architecture and (b) an autoregressive model composed of fully connected layers. Unlike existing state-of-the-art methods that operate on the entire computational domain, our CNN-based autoencoder computes a lower-dimensional basis for solution and condition fields represented on subdomains. Timestepping is performed entirely in the latent space, generating embeddings of the solution variables from the time history of embeddings of solution and condition variables. This approach not only reduces computational complexity but also enhances scalability, making it well-suited for large-scale simulations. Furthermore, to improve the stability of our rollouts, we employ a curriculum learning (CL) approach during the training of the autoregressive model. The domain-decomposition strategy enables scaling to out-of-distribution domain sizes while maintaining the accuracy of predictions – a feature not easily integrated into popular DL-based approaches for physics simulations. We benchmark our model against two widely-used DL architectures, Fourier Neural Operator (FNO) and U-Net, and demonstrate that our framework outperforms them in terms of accuracy, extrapolation to unseen timesteps, and stability for a wide range of use cases.
@@ I Introduction
In recent years, the intersection of deep learning (ML) and physics-based modeling has gained significant attention, fueled by the potential of data-driven techniques to enhance our understanding and prediction of complex physical phenomena. Several methods, both data-driven and physics-informedRaissi and Karniadakis (2018); Li et al. (2020, 2023); Lu et al. (2021); Pfaff et al. (2020); Sanchez-Gonzalez et al. (2020); Stachenfeld et al. (2021), have been developed to create surrogates for solving partial differential equations (PDEs). These deep learning (DL) models offer the promise of reducing the computational time required to solve highly nonlinear PDEs. In addition to developing purely data-driven DL models, there have been efforts to create hybrid models that combine the strengths of traditional PDE solvers with the capabilities of DL methods (Ranade et al., 2022a, b; Ranade, Hill, and Pathak, 2021; He et al., 2023; Zhang et al., 2022; Kahana et al., 2023). These hybrid approaches leverage the generalization ability of PDE solvers while using DL techniques to effectively model high-dimensional manifolds by identifying repeated patterns in the solution and geometry fields. As a result, they serve as an augmentation to traditional PDE solvers(Bar-Sinai et al., 2019; Kochkov et al., 2021; Um et al., 2020; Tathawadekar et al., 2021; Brahmachary and Thuerey, 2024; Ramos, Trost, and Thuerey, 2022; List, Chen, and Thuerey, 2022; Coros et al., 2021; Holl, Koltun, and Thuerey, 2020). These DL-based models have demonstrated significant potential in fields where PDE modeling is essential, such as weather prediction (Gupta and Brandstetter, 2022; Lam et al., 2022), 3-D aerodynamics predictions Li et al. (2024), etc.
@@
Among the various frameworks developed for physics-based problems, physics-informed neural networks (PINNs) Raissi and Karniadakis (2018) and Fourier neural operators (FNOs) Li et al. (2020, 2024) are two of the most prevalent, both with distinct strengths and applications. PINNs leverage underlying PDEs by incorporating residual governing equations directly into the loss function during training. This approach allows them to solve both forward and inverse problems while respecting physical laws, making them particularly useful in scenarios where data is scarce but the physics is well understood. In contrast, FNOs take a different approach by learning operators in the frequency domain. By using Fourier transforms, FNOs capture global information and efficiently handle complex, high-dimensional problems. In recent years, graph-based approaches using graph neural networks (GNNs) have also seen significant growth, as demonstrated by the works of Brandstetter et al. Brandstetter, Worrall, and Welling (2022), Xu et al. Xu, Pradhan, and Duraisamy (2021), and Boussif et al. Boussif et al. (2022). One of the key advantages of GNN-based methods is the seamless formulations of unstructured simulations, prevalant in PDE modeling of complex physics Mavriplis (1997), on graphs.
@@
One of the key distinctions of existing deep learning methods, such as those mentioned earlier, is that most of them are global in nature. By ‘global’, we mean that they learn the solution distribution across the entire computational domain simultaneously. This contrasts with traditional PDE solvers, which derive solutions at computational grid points based on local differentials. Due to their global learning approach, these deep learning models often struggle with accurately resolving small-scale features or extrapolating to out-of-distribution domain sizes. Moreover, they face additional challenges such as the inability to represent complex geometries, difficulty in scaling to high-dimensional problems with the high-resolution meshes that are quite common in industrial simulations, exorbitantly large training times, prohibitive GPU memory requirements, and limited generalization to out-of-distribution input spaces. These issues significantly hinder the practical application of deep learning in real-world problems. Inspired by the local approach of traditional PDE solvers, Ranade et al. Ranade et al. (2022a) introduced the Composable Machine Learning Simulator (CoMLSim) framework for steady-state PDEs. Unlike global deep learning methods, CoMLSim focuses on learning solution distributions and conditions in low-dimensional spaces within local subdomains rather than across the entire computational domain. It then stitches these local solutions together by ensuring consistency and continuity across subdomains using flux schemes learned with neural networks. This approach integrates two key features: (a) a local-learning strategy and (b) a low-dimensional learning. Ranade et al. demonstrated that CoMLSim outperformed DL methods learning on the global level, such as FNO Li et al. (2020), DeepONet Lu et al. (2021), and U-NetRonneberger, Fischer, and Brox (2015) on a range of steady-state PDEs, from simple Poisson’s equations to industrial applications like 3D Reynolds Averaged Navier-Stokes (RANS) equations and chip coolingRanade et al. (2022b).
@@ III.1 Accuracy comparison with baselines for the interpolation task
Figures 5, 6, 7 present three randomly selected contours obtained using transient-CoMLSim at last timesteps from the 2-D shallow water, 2-D vortex-flow, and 2-D diffusion-reaction, respectively. From the contours, it is evident that the transient-CoMLSim captures the physics of the respective 2-D datasets quite accurately throughout the course of their respective temporal evolution.
@@ III.2 Extrapolation to unseen timesteps and bigger domain sizes
Table 5 presents the nRMSE of the transient-CoMLSim model on the six test datasets of the 3-D additive manufacturing case. The number of timesteps differ from that used during training – $N_{t}=400$ during training for the 3-D additive manufacturing case. Additionally, datasets AM4, AM5, and AM6 also test the model’s ability to extrapolate to larger spatial domains, as discussed in the context of table 2. Table 5 demonstrates that the model shows remarkable accuracy for all six test samples, effectively extrapolating to unseen timesteps and domain sizes. Figure 8 presents the temperature contours at the top-most plane for all six test samples, at the end of their respective simulation time. The model faithfully captures the spatial distribution of the temperature in the 3-D additive manufacturing case, similar to its 2-D counterparts that are discussed in the previous section. Figure 9 shows the temporal evolution of the melt pool depth, a global quantity, for all six test samples. In the context of additive manufacturing, the melt pool depth is the depth to which the temperature goes beyond melting point during simulation. Transient-CoMLSim demonstrates remarkable accuracy in capturing the temporal trend of the melt pool depth for all six test cases.
III.3 Ablation studies for transient-CoMLSim on 2-D SWE dataset

### v3

@@ Abstract
In this paper, we propose a domain-decomposition-based deep learning (DL) framework, named transient-CoMLSim, to accurately model unstable and nonlinear partial differential equations (PDEs). The framework consists of two key components: (a) a convolutional neural network (CNN)-based autoencoder architecture and (b) an autoregressive model composed of fully connected layers. Unlike existing state-of-the-art methods that operate on the entire computational domain, our CNN-based autoencoder computes a lower-dimensional basis for solution and condition fields represented on subdomains. Timestepping is performed entirely in the latent space, generating embeddings of the solution variables from the time history of embeddings of solution and condition variables. This approach not only reduces computational complexity, but also enhances scalability, making it well-suited for large-scale simulations. Furthermore, to improve the stability of our rollouts, we employ a curriculum learning (CL) approach during the training of the autoregressive model. The domain-decomposition strategy enables scaling to out-of-distribution domain sizes while maintaining the accuracy of predictions – a feature not easily integrated into popular DL-based approaches for physics simulations. We benchmark our model against two widely used DL architectures, Fourier Neural Operator (FNO) and U-Net, and demonstrate that our framework outperforms them in terms of accuracy, extrapolation to unseen timesteps, and stability for a wide range of use cases.
@@ I Introduction
In recent years, the intersection of deep learning (ML) and physics-based modeling has gained significant attention, fueled by the potential of data-driven techniques to enhance our understanding and prediction of complex physical phenomena. Several methods, both data-driven and physics-informedRaissi and Karniadakis (2018); Li et al. (2020, 2023); Lu et al. (2021); Pfaff et al. (2020); Sanchez-Gonzalez et al. (2020); Stachenfeld et al. (2021), have been developed to create surrogates to solve partial differential equations (PDEs). These deep learning (DL) models offer the promise of reducing the computational time required to solve highly nonlinear PDEs. In addition to developing DL models driven purely by data, there have been efforts to create hybrid models that combine the strengths of traditional PDE solvers with the capabilities of DL methods (Ranade et al., 2022a, b; Ranade, Hill, and Pathak, 2021; He et al., 2023; Zhang et al., 2022; Kahana et al., 2023). These hybrid approaches leverage the generalization ability of PDE solvers while using DL techniques to effectively model high-dimensional manifolds by identifying repeated patterns in the solution and geometry fields. As a result, they serve as an augmentation of traditional PDE solvers(Bar-Sinai et al., 2019; Kochkov et al., 2021; Um et al., 2020; Tathawadekar et al., 2021; Brahmachary and Thuerey, 2024; Ramos, Trost, and Thuerey, 2022; List, Chen, and Thuerey, 2022; Coros et al., 2021; Holl, Koltun, and Thuerey, 2020). These DL-based models have demonstrated significant potential in fields where PDE modeling is essential, such as weather prediction (Gupta and Brandstetter, 2022; Lam et al., 2022), 3-D aerodynamics predictions Li et al. (2024), etc.
@@
Among the various frameworks developed for physics-based problems, physics-informed neural networks (PINNs) Raissi and Karniadakis (2018) and Fourier neural operators (FNOs) Li et al. (2020, 2024) are two of the most prevalent, both with distinct strengths and applications. PINNs leverage the underlying PDEs by incorporating the residual governing equations directly into the loss function during training. This approach allows them to solve both forward and inverse problems while respecting physical laws, making them particularly useful in scenarios where data is scarce but the physics is well understood. In contrast, FNOs take a different approach by learning operators in the frequency domain. Using Fourier transforms, FNOs capture global information and efficiently handle complex, high-dimensional problems. In recent years, graph-based approaches using graph neural networks (GNNs) have also seen significant growth, as demonstrated by the works of Brandstetter et al. Brandstetter, Worrall, and Welling (2022), Xu et al. Xu, Pradhan, and Duraisamy (2021), and Boussif et al. Boussif et al. (2022). One of the key advantages of GNN-based methods is the seamless formulation of unstructured simulations, prevalent in PDE modeling of complex physics Mavriplis (1997), on graphs.
@@
One of the key distinctions of existing deep learning methods, such as those mentioned above, is that most of them are global in nature. By ‘global’, we mean that they learn the solution distribution across the entire computational domain simultaneously. This contrasts with traditional PDE solvers, which derive solutions at computational grid points based on local differentials. Due to their global learning approach, these deep learning models often struggle to accurately resolving small-scale features or extrapolating to out-of-distribution domain sizes. Moreover, they face additional challenges such as the inability to represent complex geometries, difficulty in scaling to high-dimensional problems with the high-resolution meshes that are quite common in industrial simulations, exorbitantly large training times, prohibitive GPU memory requirements, and limited generalization to out-of-distribution input spaces. These issues significantly hinder the practical application of deep learning in real-world problems. Inspired by the local approach of traditional PDE solvers, Ranade et al. Ranade et al. (2022a) introduced the Composable Machine Learning Simulator (CoMLSim) framework for steady-state PDEs. Unlike global deep learning methods, CoMLSim focuses on learning solution distributions and conditions in low-dimensional spaces within local subdomains rather than across the entire computational domain. It then stitches these local solutions together by ensuring consistency and continuity across subdomains using flux schemes learned with neural networks. This approach integrates two key features: (a) a local-learning strategy and (b) a low-dimensional learning. Ranade et al. demonstrated that CoMLSim outperformed DL methods learning on the global level, such as FNO Li et al. (2020), DeepONet Lu et al. (2021), and U-NetRonneberger, Fischer, and Brox (2015) on a range of steady-state PDEs, from simple Poisson’s equations to industrial applications like 3D Reynolds Averaged Navier-Stokes (RANS) equations and chip coolingRanade et al. (2022b).
@@ III.1 Accuracy comparison with baselines for the interpolation task
Figures 5, 6, 7 present three randomly selected contours obtained using transient-CoMLSim at last timesteps from the 2-D shallow water, 2-D vortex-flow, and 2-D diffusion-reaction, respectively. From the contours, it is evident that the transient-CoMLSim captures the physics of the respective 2-D datasets quite accurately throughout the course of their respective temporal evolution. Appendix A presents animations of three randomly selected test samples, showing the time evolution of solution variables for the three 2-D PDE datasets (figures 1021).
@@ III.2 Extrapolation to unseen timesteps and bigger domain sizes
[…]
III.3 Ablation studies for transient-CoMLSim on vortex-flow dataset
@@ IV Conclusions
Appendix A Animations for the four datasets
[…]
Figure 10: SWE animation - random sample 1 (multimedia available online).
Figure 11: SWE animation - random sample 2 (multimedia available online).
Figure 12: SWE animation - random sample 3 (multimedia available online).
Figure 13: Vortex flow animation - random sample 1 (multimedia available online).
Figure 14: Vortex flow animation - random sample 2 (multimedia available online).
Figure 15: Vortex flow animation - random sample 3 (multimedia available online).
Figure 16: Diffusion reaction animation - random sample 1 - variable $u$ (multimedia available online).
Figure 17: Diffusion reaction animation - random sample 1 - variable $v$ (multimedia available online).
Figure 18: Diffusion reaction animation - random sample 2 - variable $u$ (multimedia available online).
Figure 19: Diffusion reaction animation - random sample 2 - variable $v$ (multimedia available online).
Figure 20: Diffusion reaction animation - random sample 3 - variable $u$ (multimedia available online).
Figure 21: Diffusion reaction animation - random sample 3 - variable $v$ (multimedia available online).
Figure 22: AM1 animation (multimedia available online).
Figure 23: AM2 animation (multimedia available online).
Figure 24: AM3 animation (multimedia available online).
Figure 25: AM4 animation (multimedia available online).
Figure 26: AM5 animation (multimedia available online).